OpenStudy (anonymous):

I give medals!!! Solve for x: 3(x + 1) = -2(x - 1) + 6

OpenStudy (anonymous):

First Distribute 3x+3=2x-2+6 3x+3=2x+4 -2x -2x x+3=4 -3 -3 x=4

OpenStudy (anonymous):

Thanks! how about this last one? Solve for x: -3(x + 3) = -3(x + 1) - 5

OpenStudy (anonymous):

Distribute again -3x-9=-3x-3-5 -3x-9=-3x-8 +3x +3x -9 does not equal -8 so no real solutions

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

are there answer choices

OpenStudy (anonymous):

6 -6 All real numbers No solution

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

looks like you're right :)

OpenStudy (anonymous):

thanks <3

OpenStudy (anonymous):

no problem bae